OpenStudy (anonymous):

Solve, using the substitution method. m + n = 5 m – n = 3 (Points : 1) There is no solution. The solution is (2, –1) The solution is (4, 1) There are an infinite number of solutions.

OpenStudy (anonymous):

(4,1)

OpenStudy (anonymous):

m + n = 5 --> m = 5 - n now sub 5 - n in for m m - n = 3 5 - n - n = 3 5 - 2n = 3 -2n = 3 - 5 -2n = -2 n = 1 now sub 1 in for n in either of the original equations to find m m + n = 5 m + 1 = 5 m = 5 - 1 m = 4 solution is : m = 4 and n = 1

OpenStudy (anonymous):

thank u for explaining @kelliegirl33

OpenStudy (anonymous):

no problem :)